# Speeding up an unsteady flow simulationM. HANEK, J. PAPEŽ, and J. ŠÍSTEK

### v1

@@ Abstract
We deal with accelerating the solution of a sequence of large linear systems solved by an iterative Krylov subspace method. The sequence originates from time-stepping within a simulation of an unsteady incompressible flow. We apply a pressure correction scheme, and we focus on the solution of the Poisson problem for the pressure corrector. Its scalable solution presents the main computational challenge in many applications. The right-hand side of the problem changes in each time step, while the system matrix is constant and symmetric positive definite. The acceleration techniques are studied on a particular problem of flow around a unit sphere. Our baseline approach is based on a parallel solution of each problem in the sequence by nonoverlapping domain decomposition method. The interface problem is solved by the preconditioned conjugate gradient (PCG) method with the three-level BDDC preconditioner. Three techniques for accelerating the solution are gradually added to the baseline approach. First, the stopping criterion for the PCG iterations is studied. Next, deflation is used within the conjugate gradient method with several approaches to Krylov subspace recycling. Finally, we add the adaptive selection of the coarse space within the three-level BDDC method. The paper is rich in experiments with careful measurements of computational times on a parallel supercomputer. The combination of the acceleration techniques eventually leads to saving about one half of the computational time.
@@ 1 Introduction
We study the problem of solving a sequence of linear systems with a constant matrix and a number of right-hand sides. There are many scenarios resulting in such sequences, and we apply and study the methods on the Poisson problem of pressure (corrector) within an unsteady simulation of an incompressible flow.
@@
Unsteady incompressible flows of Newtonian fluids are modelled by the Navier-Stokes equations. The system is time-dependent and nonlinear, and in its generality, it leads to solving a system of nonlinear equations repeatedly in each time step. This approach can be very demanding with respect to computational resources, and a number of widely used numerical schemes circumvent the need to solve these systems by splitting the whole problem into solving a sequence of simpler problems within each time step. Let us mention the Pressure Implicit with Splitting of Operator (PISO) scheme [13], implicit-explicit (IMEX) time integration [14], or the pressure-correction methods (see, e.g., the review in [10]), to name a few. In these approaches, a Poisson-type problem for pressure typically becomes the most time-consuming problem to solve, and this trend worsens with the problem size, see, e.g., [23]. This is the reason why a considerable attention has been devoted to accelerating the Poisson-type problems in this context in literature (e.g., [3, 4, 5, 9, 19]). To solve the problem in parallel, multigrid methods, preconditioned Krylov subspace methods, and combination thereof are often employed (see, e.g., [28]).
@@
Somewhat less common is using domain decomposition solvers in this context. We investigated the applicability of a nonoverlapping domain decomposition method, namely the Balancing Domain Decomposition based on Constrains (BDDC) preconditioner, for this task in [11]. It provides the baseline approach also for the present paper. More specifically, we use the three-level BDDC preconditioner [27] for the preconditioned conjugate gradient (PCG) method. Domain decomposition (DD) methods have a relatively expensive setup, which includes factorization of the local subdomain matrices, and the matrix of the coarse problem. It is then repeatedly used for each right-hand side in the time-stepping sequence. The DD method reduces a global problem to the interface among subdomains, and runs PCG on the Schur complement problem. This leads to a significant reduction of the size of the vectors in the Krylov method.
@@
We apply the finite element method (FEM) in connection with the incremental pressure-correction method to discretize the problem. If the computational domain is fixed, the problems for the pressure corrector have a constant matrix and a new right-hand side vector in each time step. Hence, the interface Schur complement problem and the BDDC preconditioner are set up once in the first time step and reused in subsequent time steps.
@@
In the rest of this section, we briefly review the techniques employed to accelerate the solution of the arising sequence of linear systems. The gradual improvements eventually led us to saving up to one half of the overall simulation time for realistic computations. The techniques are then described in more detail in the subsequent sections.
@@
When solving a sequence of algebraic systems corresponding to time-dependent simulations with subsequent solutions close one to another, it seems appropriate to use the previously computed approximation as the starting guess for the current (new) system. This however may not bring a significant improvement in the solution time unless the stopping criterion is set properly, as we will discuss and illustrate in numerical experiments.
In this study we terminate the iterations of PCG after a sufficient reduction of the relative residual. However, we suggest to normalize the residual norm $\|{\bf r}_{k}\|$ (where ${\bf r}_{k}={\bf b}-{\bf A}{\bf x}_{k}$ is the residual associated with the approximation ${\bf x}_{k}$ from the $k$th iteration) by a quantity that is not related to the initial approximation. For that purpose, we take the norm of the right-hand side ${\bf b}$ and terminate the iterations when
@@
$\frac{\|{\bf r}_{k}\|}{\|{\bf r}_{0}\|}<10^{-6}.$
We show that the latter criterion does not reduce the number of iterations in case when a good initial approximation is available.
@@
Running the conjugate gradient (CG) method in a subspace defined as the orthogonal complement of a stored basis is an established approach known as deflated CG [8, 22]. Since Krylov subspace methods typically reduce oscillating component of the error faster than smooth components, making the Krylov subspace orthogonal to the eigenvectors corresponding to the smallest eigenvalues typically results in faster convergence. As such, deflation can be seen as an alternative to preconditioning (see, e.g., [25]). On the other hand, it requires a careful implementation as some of the operations in the approach are sensitive to accumulation of rounding errors and numerical loss of orthogonality. Another drawback of these methods is the relatively high memory cost of storing the deflation basis which limits its size. Both of these issues are mitigated by the use of the BDDC preconditioner in our context.
@@
The BDDC preconditioner allows a flexible definition of its coarse space through defining the coarse degrees of freedom. There are several variants of the adaptive BDDC (see the overview papers [15, 21]), and the common feature of the methods is solving a number of local eigenvalue problems, followed by using selected eigenvectors for defining optimal coarse degrees of freedom. The enriched coarse space typically improves the convergence of the iterative method at the cost of more expensive preconditioner setup due to solving the eigenvalue problems. While these approaches were developed mostly for problems with heterogeneous materials with largely varying coefficients, they can be seen as a general approach to adjusting the strength of the preconditioner. Hence, they present an interesting option also for time-dependent problems studied in this paper, in which any reduction of the number of iterations can lead to large savings of the computational time needed for the whole sequence.
In our setting, we combine the adaptive BDDC preconditioner with deflated CG, so that we can benefit from synergy of these approaches in several aspects. Namely, the CG method runs on much shorter vectors (with the length given by the interface size rather than the global size) allowing storing more basis vectors of the deflation basis. A nonstandard feature of combining BDDC with deflated CG is that the spectrum of the preconditioned operator is bounded from below by 1 with many eigenvalues clustered near this value. This means that deflating eigenvectors corresponding to the smallest eigenvalues does not significantly improve the CG convergence. Instead, we use the eigenvectors corresponding to the largest eigenvalues of the preconditioned operator as the basis for deflation. This can be understood as another level of preconditioning, further reducing the upper bound of the spectrum of the preconditioned operator; see, e.g., [25].
In our case study (described in more details in Section 6) the proper choice of stopping criterion can save 14 % of the computational time spent on solving the system for pressure correction. When combined with a subspace deflation, this improvement is 28 %. Using adaptive coarse space with the proper stopping criterion saves 32 %. Finally, the combination of the proper stopping criterion, subspace deflation, and adaptive coarse space reduces the time by 36 %. The savings are even higher, over 50 %, in a larger tested problem.
@@ 2 Model problem
(1) $\frac{\partial\bm{u}}{\partial t}+(\bm{u}\cdot\nabla)\bm{u}-\nu\Delta\bm{u}+\nabla p$ $=\bm{0}\quad\mbox{in}\ \Omega,$
@@
where $\bm{u}$ is the velocity vector of the fluid, $t$ denotes time, $\nu$ is the kinematic viscosity of the fluid, and $p$ is the kinematic pressure. System (1) is complemented by the initial and boundary conditions: $\bm{u}(t=0,\bm{x})=\bm{0}$ in $\Omega$, $\bm{u}(t,\bm{x})=\bm{u}_{\infty}$ on $\partial\Omega_{\infty}$, $\bm{u}(t,\bm{x})=\bm{0}$ on $\partial\Omega_{S}$, and $-\nu(\nabla\bm{u}){\bm{n}}+p{\bm{n}}=\bm{0}$ on $\partial\Omega_{O}$, with $\bm{n}$ being the unit outer normal vector of $\partial\Omega$.
@@
    (2) $\frac{1}{\Delta t}\bm{u}^{(n+1)}+(\bm{u}^{(n)}\cdot\nabla)\bm{u}^{(n+1)}-\nu\Delta\bm{u}^{(n+1)}=\frac{1}{\Delta t}\bm{u}^{(n)}-\nabla(p^{(n)}+\psi^{(n)})\quad\mbox{in}\ \Omega$
@@
    (3) $-\Delta\psi^{(n+1)}=-\frac{1}{\Delta t}\nabla\cdot\bm{u}^{(n+1)}\quad\mbox{in}\ \Omega$
@@
    (4) $p^{(n+1)}=p^{(n)}+\psi^{(n+1)}-\nu\nabla\cdot\bm{u}^{(n+1)}.$
Problems (2), (3), and (4) are discretized by the finite element method (FEM) using Taylor-Hood $Q_{2}-Q_{1}$ hexahedral elements. They approximate the velocity and pressure fields by continuous piecewise tri-quadratic and tri-linear basis functions, respectively. In the finite element mesh, there are $n_{\bm{u}}$ nodes with velocity unknowns and $n_{p}$ nodes with pressure unknowns, with the ratio $n_{\bm{u}}/n_{p}$ being approximately 8.
For solving the algebraic problems arising from (2) and (4), we use the methods identified as optimal by [23]. In particular, the Generalized Minimal Residual method (GMRES) is used for (2), and the PCG method for (4). Block Jacobi preconditioner using ILU(0) on the blocks corresponding to mesh partitions is used for both problems.
@@
(5) ${\bf K}\,{\bf x}^{(n+1)}={\bf f}^{(n+1)}\,.$
This is a well-studied case from the point of view of domain decomposition methods, which are suitable solvers for this task. The main focus of this study is a scalable solution of (5) arising from the Poisson problem for pressure corrector (3).
@@ 3 BDDC method with adaptive coarse space
As the baseline approach to solve the sequence of problems (5), we employ the three-level version [27, 17] of the BDDC method [6] implemented in the BDDCML111 https://users.math.cas.cz/~sistek/software/bddcml.html library. In particular, the approach is based on (i) reducing the global problem on the whole domain to the reduced problem defined at the interface between subdomains, (ii) solving the reduced problem using PCG, while (iii) preconditioning the problem by the BDDC preconditioner. For the sake of brevity, we drop the time index $(n+1)$ in the following discussion. More precisely, vectors ${\bf x}$ and ${\bf f}$ without superscript are considered in the $(n+1)$-st time level, while the superscript $(n)$ will be kept for their counterparts from the previous $n$-th time level.
@@ 3.1 Iterative substructuring
First, we consider the reduction of the global problem to the inter-subdomain interface. This procedure is rather standard and described, e.g., in monographs [7, 26]. To this end, the finite element mesh is divided into $N_{S}$ nonoverlapping subdomains $\Omega_{i},i=1,\dots,N_{S}$, with the partition respecting inter-element boundaries. The subset of unknowns to which elements of subdomain $\Omega_{i}$ contribute to is called local subdomain unknowns. This construction splits the unknowns to those belonging to just one subdomain, called interior unknowns, and the unknowns shared by several subdomains, which form the interface $\Gamma$.
The first step in iterative substructuring is to reduce the global problem to the interface. For each local subdomain, we split the unknowns belonging to interior ${\bf x}_{i}^{I}$ and the interface ${\bf x}_{i}^{\Gamma}$, which leads to a 2$\times$2 blocking of the local linear system,
(6) $\left[\begin{array}[c]{cc}{\bf K}_{i}^{II}&{\bf K}_{i}^{I\Gamma}\\ {\bf K}_{i}^{\Gamma I}&{\bf K}_{i}^{\Gamma\Gamma}\end{array}\right]\left[\begin{array}[c]{cc}{\bf x}_{i}^{I}\\ {\bf x}_{i}^{\Gamma}\end{array}\right]=\left[\begin{array}[c]{cc}{\bf f}_{i}^{I}\\ {\bf f}_{i}^{\Gamma}\end{array}\right]\,.$
Here the local interface unknowns ${\bf x}_{i}^{\Gamma}$ can be assembled into a global set of unknowns at the interface of all subdomains $\Gamma$ utilising an interface restriction matrix ${\bf R}^{\Gamma}_{i}:\Gamma\rightarrow\Gamma_{i}$, where ${\bf x}^{\Gamma}=\sum_{i=1}^{N_{S}}\left({\bf R}_{i}^{\Gamma}\right)^{T}{\bf x}_{i}^{\Gamma}$.
The iterative substructuring seeks the solution of the global interface problem
(7) ${\bf A}{\bf x}^{\Gamma}={\bf b}\,,$
using, for instance, the PCG method. The global Schur complement matrix ${\bf A}$ comprises of the subdomain contributions,
(8) ${\bf A}=\sum_{i=1}^{N_{S}}\left({\bf R}_{i}^{\Gamma}\right)^{T}{\bf A}_{i}{\bf R}^{\Gamma}_{i}\,,$
where the local Schur complement with respect to $\Gamma_{i}$ is defined as
(9) ${\bf A}_{i}={\bf K}_{i}^{\Gamma\Gamma}-{\bf K}_{i}^{\Gamma I}\left({\bf K}_{i}^{II}\right)^{-1}{\bf K}_{i}^{I\Gamma}\,.$
Similarly, the right-hand side is assembled from the subdomain vectors
(10) ${\bf b}=\sum_{i=1}^{N_{S}}\left({\bf R}_{i}^{\Gamma}\right)^{T}{\bf b}_{i}\,,$
where
(11) ${\bf b}_{i}={\bf f}_{i}^{\Gamma}-{\bf K}_{i}^{\Gamma I}\left({\bf K}_{i}^{II}\right)^{-1}{\bf f}_{i}^{I}.$
Once we know the local solution at the interface ${\bf x}_{i}^{\Gamma}$, the solution in the interior of each subdomain ${\bf x}_{i}^{I}$ is recovered from the first row of (6). Note that neither the global Schur complement matrix ${\bf A}$ nor the local Schur complement matrices ${\bf A}_{i}$ are explicitly constructed in the iterative substructuring, since only multiplications of vectors by matrices ${\bf A}_{i}$ are needed at each PCG iteration.

### v3

# Speeding up an unsteady flow simulation by adaptive BDDC and Krylov subspace recycling

@@ Abstract
We deal with accelerating the solution of a sequence of large linear systems solved by preconditioned conjugate gradient method (PCG). The sequence originates from time-stepping within a simulation of an unsteady incompressible flow. We apply a pressure correction scheme and focus on the solution of the Poisson problem for the pressure corrector. Its scalable solution presents the main computational challenge in many applications. The right-hand side of the problem changes in each time step, while the system matrix is constant and symmetric positive definite. The acceleration techniques are studied on a representative problem of flow around a unit sphere. Our baseline approach is based on a parallel solution of each problem in the sequence by nonoverlapping domain decomposition method. The interface problem is solved by PCG with the three-level BDDC preconditioner. As a preliminary step, an appropriate stopping criterion for the PCG iterations is chosen. Next, two techniques for accelerating the solution are gradually added to the baseline approach. Deflation is used within PCG with several approaches to Krylov subspace recycling. Finally, we add the adaptive selection of the coarse space within the three-level BDDC method. The paper is rich in experiments with careful measurements of computational times on a parallel supercomputer. The combination of the acceleration techniques eventually leads to saving more than 40 % of the computational time.
@@ 1 Introduction
We study the problem of solving a sequence of linear systems with a constant matrix and variable right-hand sides. There are many scenarios resulting in such sequences, and we apply and study the methods on the Poisson problem of pressure (corrector) within an unsteady simulation of an incompressible flow. As a representative model problem, we consider an unsteady flow around a unit sphere.
@@
Unsteady incompressible flows of Newtonian fluids are modelled by the Navier-Stokes equations. The system is time-dependent and nonlinear, and in its generality, it leads to solving a system of nonlinear equations repeatedly in each time step. This approach can be very demanding with respect to computational resources, and a number of widely used numerical schemes circumvent the need to solve these systems by splitting the whole problem into solving a sequence of simpler problems within each time step. Let us mention the Pressure Implicit with Splitting of Operator (PISO) scheme [1], implicit-explicit (IMEX) time integration [2], or the pressure-correction methods (see, e.g., the review in [3]), to name a few. In these approaches, a Poisson-type problem for pressure often becomes the most time-consuming problem to solve, and this trend worsens with the problem size, see, e.g., [4]. This is the reason why a considerable attention has been devoted in literature to accelerating the Poisson-type problems in this context (e.g., [5, 6, 7, 8, 9]). To solve the problem in parallel, multigrid methods, preconditioned Krylov subspace methods, and combination thereof are often employed (see, e.g., [10]).
@@
While using domain decomposition solvers for stand-alone Poisson problems is well established and studied in literature, their application in the context of sequences of Poisson problems arising in incompressible flow simulations is much less common. We investigated the applicability of a nonoverlapping domain decomposition method for this task in [11], considering the Balancing Domain Decomposition based on Constrains (BDDC) preconditioner introduced in [12]. It provides the baseline approach also for the present paper. More specifically, we use the three-level BDDC preconditioner [13, 14] for the preconditioned conjugate gradient (PCG) method. Domain decomposition (DD) methods have a relatively expensive setup, which includes factorization of the local subdomain matrices, and the matrix of the coarse problem. It is then repeatedly used for each right-hand side in the time-stepping sequence. The DD method reduces a global problem to the interface among subdomains, and runs PCG on the Schur complement problem. This leads to a significant reduction of the size of the vectors in the Krylov method.
@@
We apply the finite element method (FEM) in connection with the incremental pressure-correction scheme to discretize the problem. If the computational domain is fixed, the problems for the pressure corrector have a constant matrix and a new right-hand side vector in each time step. Hence, the interface Schur complement problem and the BDDC preconditioner are set up only in the first time step and reused in the subsequent time steps.
@@
In the rest of this section, we briefly review the techniques employed to accelerate the solution of the arising sequence of linear systems. The gradual improvements eventually led us to saving more than 40 percent of the overall simulation time for realistic computations. The techniques are then described in more detail in the subsequent sections.
@@
When solving a sequence of algebraic systems corresponding to time-dependent simulations with subsequent solutions close one to another, it seems appropriate to use the previously computed approximation as the starting guess for the current (new) system. This however may not bring a significant improvement in the solution time unless an appropriate stopping criterion is used, as we will discuss and illustrate in numerical experiments.
In this study we terminate the iterations of PCG after a sufficient reduction of the relative residual. However, we suggest to normalize the residual norm $\|{\bf r}_{k}\|$ (where ${\bf r}_{k}={\bf b}-{\bf A}{\bf x}_{k}$ is the residual associated with the approximation ${\bf x}_{k}$ from the $k$-th iteration) by a quantity that is not related to the initial approximation. For that purpose, we take the norm of the right-hand-side vector $\bf b$} and terminate the iterations when
@@
$\frac{\|{\bf r}_{k}\|}{\|{\bf r}_{0}\|}<10^{-6},$
which is inappropriate in the cases when a good initial approximation ${\bf x}_{0}$ is available.
@@
This technique, which involves reusing information from previous runs of a Krylov method, is frequently employed for nonsymmetric systems and restarted methods; see, e.g., [15] or the survey [16]. Applying a subspace recycling to improve the convergence of the conjugate gradient (CG) method for solving a sequence of SPD linear systems has been considered, e.g., in [17, 18]. Running the CG method in a subspace defined as the orthogonal complement of a stored basis is an established approach known as deflated CG [19, 20, 17]. Since CG typically reduces oscillating components of the error faster than smooth components, making the Krylov subspace orthogonal to the eigenvectors corresponding to the smallest eigenvalues typically results in faster convergence. As such, deflation can be seen as an alternative to preconditioning; see, e.g., [21] and [22]. The deflation, on the other hand, requires a careful implementation as some of the operations in the approach are sensitive to accumulation of rounding errors and numerical loss of orthogonality. Another drawback of these methods might be the relatively high memory cost of storing the deflation basis which limits its size. Both of these issues are mitigated by the use of the BDDC preconditioner in our context. This synergy was already realized for the earlier Finite Element Tearing and Interconnecting (FETI) method [23] applied to sequences of right-hand sides in [24]. In [22], the deflation is seen as an alternative way to implementing the coarse space in BDDC and FETI-DP methods.
@@
The BDDC preconditioner allows a flexible definition of its coarse space through defining the coarse degrees of freedom. In particular, there are several variants of the adaptive BDDC (see the overview papers [25, 26]), and the common feature of the methods is solving a number of local eigenvalue problems, followed by using selected eigenvectors for defining optimal coarse degrees of freedom. The enriched coarse space typically improves the convergence of the iterative method at the cost of more expensive preconditioner setup due to solving the eigenvalue problems. While these approaches were developed mostly for problems with heterogeneous materials with largely varying coefficients, they can be seen as a general approach to adjusting the strength of the DD preconditioners. For this reason, they present an interesting option also for time-dependent problems studied in this paper, in which any reduction of the number of iterations can lead to large savings of the computational time needed for the whole sequence. In the present work, we apply the method described in [27, 28]. It is combined with the three-level BDDC preconditioner as a special case of adaptive multilevel BDDC [29].
In our setting, we combine the adaptive BDDC preconditioner with deflated CG, so that we can benefit from synergy of these approaches in several aspects. Namely, the CG method runs on much shorter vectors (with the length given by the interface size rather than the global size) allowing storing more basis vectors of the deflation basis. A nonstandard feature of combining BDDC with deflated CG is that the spectrum of the preconditioned operator is bounded from below by 1 with many eigenvalues clustered near this value. This means that deflating eigenvectors corresponding to the smallest eigenvalues does not significantly improve the CG convergence. Instead, we use the eigenvectors corresponding to the largest eigenvalues of the preconditioned operator as the basis for deflation. This can be understood as another level of preconditioning, further reducing the upper bound of the spectrum of the preconditioned operator; see, e.g., [21, 22].
In our case study (described in more details in Section 6), the proper choice of stopping criterion can save 11 % of the computational time spent on solving the system for pressure correction. Considering the process with proper initial guess and stopping criterion as a baseline approach, the improvement of the subspace deflation is around 8 %. Using adaptive coarse space in BDDC preconditioner saves nearly 19 %. Finally, the combination of the subspace deflation and adaptive coarse space reduces the time by 23 %. The savings are even higher, around 43 %, for a larger tested problem. Nevertheless, the objective of this paper is not to demonstrate computational savings in a specific problem instance. It rather aims at presenting a set of techniques which, either individually or in combination, can be applied efficiently across a broad class of problems in incompressible flows.
@@ 2 Model problem
$\frac{\partial\bm{u}}{\partial t}+(\bm{u}\cdot\nabla)\bm{u}-\nu\Delta\bm{u}+\nabla p$ $=\bm{0}\quad\mbox{in}\ \Omega,$ (1)
@@
where $\bm{u}(t,\bm{x})$ is the velocity vector of the fluid, $t$ denotes time, $\nu$ is the kinematic viscosity of the fluid, and $p$ is the kinematic pressure. System (1) is complemented by the initial and boundary conditions: $\bm{u}(t=0,\bm{x})=\bm{0}$ in $\Omega$, $\bm{u}(t,\bm{x})=\bm{u}_{\infty}$ on $\partial\Omega_{\infty}$, $\bm{u}(t,\bm{x})=\bm{0}$ on $\partial\Omega_{S}$, and $\nu(\nabla\bm{u}){\bm{n}}+p{\bm{n}}=\bm{0}$ on $\partial\Omega_{O}$, with $\bm{n}$ being the unit outer normal vector of $\partial\Omega$.
@@
    $\frac{1}{\Delta t}\bm{u}^{(n+1)}+(\bm{u}^{(n)}\cdot\nabla)\bm{u}^{(n+1)}-\nu\Delta\bm{u}^{(n+1)}=\frac{1}{\Delta t}\bm{u}^{(n)}-\nabla(p^{(n)}+\psi^{(n)})\quad\mbox{in}\ \Omega$ (2)
@@
    $\Delta\psi^{(n+1)}=-\frac{1}{\Delta t}\nabla\cdot\bm{u}^{(n+1)}\quad\mbox{in}\ \Omega$ (3)
@@
    $p^{(n+1)}=p^{(n)}+\psi^{(n+1)}-\nu\nabla\cdot\bm{u}^{(n+1)}.$ (4)
Problems (2), (3), and (4) are discretized by the finite element method (FEM) using Taylor-Hood $Q_{2}-Q_{1}$ hexahedral elements. They approximate the velocity and pressure fields by continuous piecewise tri-quadratic and tri-linear basis functions, respectively. In the finite element mesh, there are $n_{\bm{u}}$ nodes with velocity unknowns and $n_{p}$ nodes with pressure unknowns, with the ratio $n_{\bm{u}}/n_{p}$ being approximately 8. Note that applying FEM to equation (4) leads to solving an algebraic system corresponding to an $L_{2}$ projection. The term $\nu\nabla\cdot\bm{u}^{(n+1)}$ on the right-hand side is discontinuous from one element to another, and hence it cannot be simply added to the pressure function.
For solving the algebraic problems arising from (2) and (4), we use the methods identified as optimal by [4]. In particular, the Generalized Minimal Residual method (GMRES) is used for (2), and the PCG method for (4), which corresponds to a linear solve related to an $L_{2}$-projection when FEM is applied. Block Jacobi preconditioner using ILU(0) on the blocks corresponding to mesh partitions is used for both problems.
@@
${\bf K}\,{\bf x}^{(n+1)}={\bf f}^{(n+1)}\,.$ (5)
This is a well-studied case from the point of view of domain decomposition methods, which are suitable solvers for this task. The main focus of this study is a scalable solution of sequence (5) arising from the Poisson problem for pressure corrector (3).
In the case of $\partial\Omega_{O}=\emptyset$, problem (3) corresponds to solving the so-called pure Neumann problem, and matrix $\bf K$} is only semidefinite. It has a one-dimensional nullspace of constant vectors corresponding to constant functions in the finite element space. The deflated CG method can be efficiently employed to cope with this issue, by adding the constant vector into the deflation basis.
@@ 3 BDDC method with adaptive coarse space
As the baseline approach to solve the sequence of problems (5), we employ the three-level version [13, 14] of the BDDC method [12] implemented in the BDDCML [^1] https://users.math.cas.cz/~sistek/software/bddcml.html library. In particular, the approach is based on (i) reducing the global problem on the whole domain to the reduced problem defined at the interface between subdomains, (ii) solving the reduced problem using PCG, while (iii) preconditioning the problem by the BDDC preconditioner. For the sake of brevity, we drop the time index $(n+1)$ in the following discussion. More precisely, vectors $\bf x$} and $\bf f$} without superscript are considered in the $(n+1)$-st time level, while the superscript $(n)$ will be kept for their counterparts from the previous $n$-th time level.
@@ 3.1 Iterative substructuring
First, we consider the reduction of the global problem to the inter-subdomain interface. This procedure is rather standard and described, e.g., in monographs [30, 31]. To this end, the finite element mesh is divided into $N_{S}$ nonoverlapping subdomains $\Omega_{i},i=1,\dots,N_{S}$, with the partition respecting inter-element boundaries. The subset of unknowns to which elements of subdomain $\Omega_{i}$ contribute is called local subdomain unknowns.
In iterative substructuring, local subdomain unknowns are further split to those belonging to just one subdomain, called interior unknowns, and the unknowns shared by several subdomains, which form the interface $\Gamma$. Then we seek the solution of the global interface problem
${\bf A}{\bf x}^{\Gamma}={\bf b}$ (6)
using, for instance, the PCG method. Here $\bf A$} is the global Schur complement of the interior unknowns, ${\bf x}^{\Gamma}$ is the part of the solution vector corresponding to the interface, and $\bf b$} is the reduced right-hand side vector. Note that the global Schur complement matrix $\bf A$} is not explicitly constructed in iterative substructuring, since only multiplications of vectors by $\bf A$} are needed at each PCG iteration.